solve by the subsitution method 6x-4y= -56
8x+79=y what is the solution set. I came up with (-4,6) am I close

Answer by Alan3354(69443)

solve by the subsitution method
6x-4y= -56
8x+79=y
---------
y = 8x+79
Sub that into the other eqn
6x-4y= -56
6x - 4(8x+79) = -56
6x - 32x-316 = -56
-26x = 260
x = -10
-------
y = -1
------
--> (-10,1)
------------
what is the solution set. I came up with (-4,6) am I close
You're on the same coordinate set.